# On the gravitational stratification of multi-fluid-multi-species plasma

### v1

@@ 7 Conclusions and outlook
The primary goal of this work is to provide gravitationally stratified initial/background fields for MFMS modeling. This goal is achieved using a simple numerical integration routine based on relaxed equilibrium assumptions. Specifically, as an initial static condition is given, we may assume that all fluids can be coupled by the collisional interactions, which may then contribute to the hydrostatic equilibrium (viz. the cHE stratification) and allow us to account for any expected ionization fractions, especially including the SE ionization fractions. Without such initial fields, it is difficult to include both gravitational stratification and ionization/recombination simultaneously, since instability or disturbances caused by initial inequilibria in the pHE stratification can lead to significant errors and even sabotage numerical modeling. The present work is even more critical when more species are needed, whose pressure scale heights are vastly different, in which case the purely hydrostatic stratification (viz. the pHE stratification) is far from being in ionization equilibrium and heavy elements effectively disappear. More specifically, the cHE stratification provides a gravitational settling for chemical fractionation, in which heavy elements are (partly) supported by collisional interactions, with no other dynamic driving mechanisms included.
The present cHE stratification also clearly shows the drift velocities accompanied by strong pressure gradients, which are more prominent in the TR. Such drift velocities (decoupling between fluids) are deemed physical, as the pressure scale heights of different fluids are more significantly different in this region. In contrast, the pHE stratification does not recover such effects, regardless of the pressure gradients, and thus there is no collisional interaction between fluids, which is a strong hypothesis.
@@
In addition, we have artificially increased (or decreased) the collision rate producing a strongly (or weakly) coupled models behaving as single fluid (or independent fluids). This parametric study briefly visualizes the importance of properly describing the interactions between fluids, which may substantially change the properties of the plasma (Hansteen et al., 1997). Moreover, the result also further addresses that Ebysus code is capable when the interactions are sufficient to couple all the fluids and returns single fluid scenario.
@@ Acknowledgements.
FZ would like to thank Nicolas Poirier, Mats Carlsson, Ben Snow for the helpful discussion, and Mikolaj Szydlarski for the technical support. This research is supported by the Research Council of Norway through its Centres of Excellence scheme, project number 262622. JMS acknowledges support by NASA contract NNG09FA40C (IRIS) and 80GSFC21C0011 (MUSE), NASA grant 80NSSC26K0018, NSF grants AGS2532363 and AGS2532187. This work used the CHIANTI atomic database. CHIANTI is a collaborative project involving George Mason University, the University of Michigan (USA), the University of Cambridge (UK), and NASA Goddard Space Flight Center (USA). Figures in this work were produced using the python package Matplotlib (Hunter, 2007). Data processing was performed using the NumPy library (Harris et al., 2020) and in-house packages EbysusRunTools_py, atom_py and Helita.
@@ Appendix A Effects of magnetic field
where $\mathbf{B}_{0}$ and $r_{0}$ are free parameters, to describe the magnetic pressure from the bottom of the photosphere. With Eq. (15), a smaller $r_{0}$ means that the magnetic field is more concentrated in the photosphere. The pressure gradient can be analytically calculated based on Eq. (15), and directly added into the numerical integration routine introduced in Section 3. With a constant $r_{0}$ and varying $\mathbf{B}_{0}$, we have obtained a series of stratifications with the SE ionization, as shown in Figure 9. Note that only the total densities of all fluids are shown for simplicity. We first acknowledge that without magnetic pressure (gradient), the cHE stratification has a lower density than the averaged 2D solution. In comparison, magnetic pressure gradients significantly alter the solution, indicating that stronger gradients support higher density. More importantly, the results demonstrate that the current numerical integration routine can be used to calculate cHE stratifications under varying magnetic fields.
Figure 9: cHE stratifications calculated with analytical magnetic pressure distributions, $r_{0}=1$ Mm, and SE ionization fractions. NEQ ionization fractions can also be used to calculate such stratifications. The total density when $B_{0}=0$ G is the same as the SE result in Figure 4.
@@
in which $L$ is the radius of the loop. This function is essentially the same as what was used by Karampelas et al. (2019). As $z$ is the height from the footpoint, the gravitational acceleration along the loop axis becomes zero at the loop apex ($z=L$). Plasma along a realistic loop is supported by plasma gas pressure and magnetic pressure/tension, and we can already see that the magnetic pressure gradient effectively shapes the 1D stratification. Particularly, Figure 10 shows that the magnetic pressure description given by Eq. (15) results in sufficient magnetic pressure gradients to support the plasma in the upper chromosphere and above, where plasma $\beta$ is typically smaller than 1. Consequently, including Eq. (16) does not significantly change the stratifications, and the corresponding results are not shown for simplicity.
Again, we note that one may directly extract needed variables along open or closed field lines (or other regions of interest) from 2D/3D solar atmospheric models, and cHE stratifications can be constructed based on the extracted variables.

### v2

@@ 7 Conclusions and outlook
The primary goal of this work is to provide gravitationally stratified initial/background fields for MFMS modeling. This goal is achieved using a simple numerical integration routine based on relaxed equilibrium assumptions. Specifically, as an initial static condition is given, we may assume that all fluids can be coupled by the collisional interactions, which may then contribute to the hydrostatic equilibrium (viz., the cHE stratification) and allow us to account for any expected ionization fractions, especially including the SE ionization fractions. Without such initial fields, it is difficult to include both gravitational stratification and ionization/recombination simultaneously, since instability or disturbances caused by initial inequilibria in the pHE stratification can lead to significant errors and even sabotage numerical modeling. The present work is even more critical when more species are needed, whose pressure scale heights are vastly different, in which case the purely hydrostatic stratification (viz., the pHE stratification) is far from being in ionization equilibrium, and heavy elements effectively disappear. The cHE stratification provides a gravitational settling for chemical fractionation, in which heavy elements are (partly) supported by collisional interactions, with no other dynamic driving mechanisms included.
The present cHE stratification also clearly shows the drift velocities accompanied by strong pressure gradients, which are more prominent in the TR. Such drift velocities are deemed physical, as the pressure scale heights of different fluids are more significantly different in this region. In contrast, the pHE stratification does not recover such effects, regardless of the pressure gradients, and thus there is no collisional interaction between fluids, which is a strong hypothesis.
@@
In addition, we have artificially increased (or decreased) the collision rate, producing strongly (or weakly) coupled models behaving as a single fluid (or independent fluids). This parametric study briefly illustrates the importance of accurately describing fluid interactions, which can substantially alter plasma properties (Hansteen et al., 1997). Moreover, the result further shows that the Ebysus code is capable of coupling all the fluids when interactions are sufficiently strong.
@@ Acknowledgements.
FZ thanks Nicolas Poirier, Mats Carlsson, Boris Gudiksen, and Ben Snow for the helpful discussion and Mikolaj Szydlarski for technical support. This research has been supported by the Research Council of Norway through its Centres of Excellence scheme, project number 262622, and by computational resources provided by Sigma2, the National Infrastructure for High Performance Computing and Data Storage in Norway. JMS acknowledges support by NASA contracts NNG09FA40C (IRIS) and 80GSFC21C0011 (MUSE), NASA grant 80NSSC26K0018, NSF grants AGS2532363 and AGS2532187. This work used the CHIANTI atomic database. CHIANTI is a collaborative project involving George Mason University, the University of Michigan (USA), the University of Cambridge (UK), and NASA Goddard Space Flight Center (USA). Figures in this work were produced using the Python package Matplotlib (Hunter, 2007). Data processing was performed using the NumPy library (Harris et al., 2020) and in-house packages EbysusRunTools_py, atom_py and Helita.
@@ Appendix A Effects of magnetic field
where $\mathbf{B}_{0}$ and $r_{0}$ are free parameters, to describe the magnetic pressure from the bottom of the photosphere. With Eq. (15), a smaller $r_{0}$ means that the magnetic field is more concentrated in the photosphere. The pressure gradient can be analytically calculated based on Eq. (15), and directly added to the numerical integration routine introduced in Section 3. With a constant $r_{0}$ and varying $\mathbf{B}_{0}$, we have obtained a series of stratifications with the SE ionization, as shown in Figure 10. Note that only the total densities of all fluids are shown for simplicity. We first acknowledge that without magnetic pressure (gradient), the cHE stratification has a lower density than the averaged 2D solution. In comparison, magnetic pressure gradients significantly alter the solution, as stronger gradients support higher density. More importantly, the results demonstrate that the current numerical integration routine can be used to calculate cHE stratifications under varying magnetic fields.
Figure 10: cHE stratifications calculated with analytical magnetic pressure distributions, $r_{0}=1$ Mm, and SE ionization fractions. Alternatively, NEQ ionization fractions can also be used. The total density when $B_{0}=0$ G is the same as the SE result in Figure 4.
@@
in which $L$ is the radius of the loop. This function is essentially the same as what was used by Karampelas et al. (2019). As $z$ is the height from the footpoint, the gravitational acceleration along the loop axis becomes zero at the loop apex ($z=L$). Plasma along a realistic loop is supported by plasma gas pressure and magnetic pressure/tension, and we can already see that the magnetic pressure gradient effectively shapes the 1D stratification. Particularly, Figure 11 shows that the magnetic pressure description given by Eq. (15) results in sufficient magnetic pressure gradients to support the plasma in the upper chromosphere and above. Consequently, including Eq. (16) does not significantly change the stratifications, and the corresponding results are thus not shown.
We note that one can directly extract the required variables along open or closed field lines (or other regions of interest) from 2D/3D solar atmospheric models, and cHE stratifications can be constructed based on these variables.